OpenStudy (anonymous):

Find the product of (3x − 7y)2. A. 9x^2 − 42xy + 49y^2 B. 9x^2 + 42xy + 49y^2 C. 9x^2 − 49y^2 D. 9x^2 + 49y^2 I got A for this, is that correct?

OpenStudy (anonymous):

(3x - 7y)² = (3x)² - 2(3x)(7y) + (7y)² => 9x² - 42xy + 49y² Yep, it's A.

OpenStudy (anonymous):

Awesome thanks! I have two more that I need checked, if you don't mind?

OpenStudy (anonymous):

Anytime. Just post it

OpenStudy (anonymous):

Find the product of (3x + 7y)^2. A. 9x^2 − 42xy + 49y^2 B. 9x^2 − 49y^2 C. 9x^2 + 49y^2 D. 9x^2 + 42xy + 49y^2 I got D for this one 7. Find the product of (3x + 7y)(3x − 7y). A. 9x^2 − 42xy + 49y^2 B. 9x^2 + 42xy + 49y^2 C. 9x^2 − 49y^2 D. 9x^2 + 49y^2 and C for this one, are those right?

OpenStudy (anonymous):

(3x + 7y)² = (3x)² + 2(3x)(7y) + (7y)² => 9x² + 42xy + 49y². That's right, D. (3x + 7y)(3x − 7y) = (3x)² - (7y)² => 9x² - 49y² Those are all correct.

OpenStudy (anonymous):

yay!! I'm getting better!! :) I do have one more but its a word problem. Would you mind reading it over?

OpenStudy (anonymous):

Shoot

OpenStudy (anonymous):

just give me one sec...

OpenStudy (anonymous):

Find the two products below. Compare and contrast, in complete sentences, the similarities and differences of the two. (x + 4)(x − 4) and (x + 4)(x + 4) (x + 4)(x − 4) and (x + 4)(x + 4) The product of (x + 4)(x - 4) is x^2 - 16 The product of (x + 4)(x + 4) is x^2 + 8x + 16 The similarities are: Its the same type of operation Both equations have the same variables and numbers The differences are: In the first equation, there is a subtraction sign instead of an addition sign like in the other problem The solution is different

OpenStudy (anonymous):

In complete sentences: The similarities between these two problem are that they have the same type of operation. They also both have the same variables and numbers in the equations. The differences would come out to be the fact that in the first equation, there is a subtraction sign instead of an addition sign like in the other problem. Also,both problem have a different product.

OpenStudy (anonymous):

That's correct. You may also mention that the first one is the difference between two squares and the second one is a perfect square.

OpenStudy (anonymous):

Okay, thanks for checking those, I really appreciate it! :)

OpenStudy (anonymous):

Welcome :D

OpenStudy (anonymous):

Welcome :D

OpenStudy (anonymous):

Thanks again :)

OpenStudy (anonymous):

I hope you have an awesome afternoon.